OpenStudy (anonymous):

Please help will give medal!! Simplify -(16)1/2. Select one: a. -8 b. 8 c. -4 d. 4

OpenStudy (anonymous):

Is this

OpenStudy (anonymous):

-16 times 1/2?

OpenStudy (anonymous):

yeah. i think so(:

OpenStudy (anonymous):

So then you're multiplying two fractions! Any whole number is itself over 1. -16 = -16/1 When you multiply two fractions, you multiply across. For example: 2/3 x 2/3 = 4/9

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

Yes ma'am. :)

OpenStudy (anonymous):

THANKS(;